OpenStudy (anonymous):

What are the intercepts of y = 4x − 2?

OpenStudy (anonymous):

y-intercept: y = 4(0) - 2 = -2 x-intercept: 0 = 4x - 2 = 1/2

OpenStudy (anonymous):

i keep getting (0.5, 0) and (0, -2)

OpenStudy (anonymous):

That should be correct

OpenStudy (anonymous):

It is,thank you.

OpenStudy (anonymous):

Glad to be of assistance